5x-3y=-14 and 2x+5y=-18
Answer by satyareddy22(84)

5x-3y=-14-----------(1)
2x+5y=-18-----------(2)
multiplying (1) by 5 & (2) by 3
25x-15y=-70
6x+15y=-54
------------
31x+0=-124 =>31x=-124=>=-124/31=-4=>x=-4
now substituting x=-4 in (2)
2(-4)+5y=-18
-8+5y=-18
5y=-18+8=-10=>5y=-10
y=-10/5=-2
y=-2